# How glass breaks – Damage explains the difference between surface and fracture energies in amorphous silica

### v3

@@ Abstract
For an elastic deformation, we expect the potential energy $\psi_{\rm{pot}}$ to equal the elastic energy $\psi_{\rm{\epsilon}}$ defined by Eq. 1. Structural evolution however, which is typically not reached through elastic deformation, will be reflected in a variation in $\psi_{\rm{pot}}$. We therefore introduce the difference denoted
$\psi_{\rm{FSE}}=\psi_{\rm{pot}}-\psi_{\rm{\epsilon}}.$ (5)
Fig. 2a shows the distribution of $\psi_{\rm{FSE}}$ along the normal to the surface for the dissected sample. When the coarse-graining is omitted and the energy difference is instead calculated by averaging over 1 Å slices, a pronounced localized peak emerges, as expected for simple surface relaxation. After coarse-graining, the profile width of $\psi_{\rm{SFE}}$ is observed to be 8 Å, reflecting the convolution of the highly localized peak with the width $w$.
@@
A coarse-grained profile for the cracked samples, taken along the line indicated in (b), is also shown. The two coarse-grained $\psi_{\rm{FSE}}$ distributions are similar. This observation suggests that the $\psi_{\rm{FSE}}$ calculated in the fractured sample is the energy change due to the formation of the free surface. It is primarily the result of the surface relaxation due to the loss of atomic connectivity and differs from the damage process shown previously that penetrates significantly deeper into the sample. This is why we refer to this quantity as the free surface energy (FSE) density. Fig. 2b depicts the free surface energy during crack propagation.
@@
Since the energy localization is consistent across different sample conditions, it appears that the free surface energy $2\gamma=\int_{-\infty}^{+\infty}\psi_{\rm{FSE}}\ dy$ should also be an effective measure for assessing the area of newly formed cracks. Fig. 2c displays crack length calculated from the global sample free surface energy $\Psi_{\rm{FSE}}$ divided by $2\gamma$. It is compared with the crack length determined by the FEMU based on the damage approach eq. (4). The two measures are very consistent although the FSE length is smaller than the FEMU length. This agreement between the two methods validates the use of $\psi_{\rm{FSE}}$ as a reliable metric for quantifying crack propagation, providing an alternative to more traditional methods and offering precise measurements that align well with continuum-based models.
@@
Furthermore, analysis of the local structure (see Supplementary Material) revealed that diffuse damage is primarily associated with changes in the ring structure, whereas variations in free surface energy correlate with changes in the coordination number of silicon atoms.
@@
Global energy equilibrium.—As noted in the introduction, experimental measurements have shown that the free surface energy resulting from surface relaxation accounts for only approximately 20% of the energy required to open a crack. This section casts the free surface energy and the damage energy we have calculated at the atomic scale into a global energy balance to identify the origin of this discrepancy.
@@
Fig. 3a illustrates the various energy contributions as a function of loading. The external work, $W_{\rm{ext}}$, is partitioned into several components: the elastic strain energy, $\Psi_{\rm{\epsilon}}$, and the free surface energy (FSE), $\Psi_{\rm{FSE}}$, whose sum is represented as $\Psi_{\rm{pot}}$. Additionally, there is a small non-linear contribution, $\Psi_{\rm{etc}}$, and a significant unquantified remainder. The non-linear energy contribution is determined by isolating the portion of the elastic energy density that exceeds the local potential energy, which ideally remains positive under the assumption of linear elasticity. This excess energy contribution, is then subtracted from the total elastic energy. However, it is noteworthy that this contribution (represented by the dashed line) is minor. This observation suggests that energy dissipation due to shear plasticity, densification or ring folding is indeed negligible. Under tensile loading followed by fracture, these mechanisms contribute only marginally to the overall energy dissipation within the material.
@@
Finally, Fig. 3c presents the fracture toughness ($g_{c}$) as a function of loading, calculated using three distinct methods: the J integral based on the coarse-grained stress and displacement fields (red filled circles – detailed calculations provided in the Supplementary Materials), the FEMU phase-field fit (open circles) and the energy balance approach ($\Psi_{d}/\Gamma_{\rm{SFE}}$) obtained from the coarse-grained atomic scale simulations. The initial linear increase follows from the crack initiation process. Subsequently, the fracture surface energy remains consistent across the three calculation methods. The resulting value provides a ratio of $2\gamma/g_{c}\approx 23$%: this value found at the atomic scale is consistent with experimental findings [11].
@@
Figure 3: Energy balance of the cracked sample as a function the loading parameter  [^3] We note that $K_{I}$ loses its precise physical significance as the tip advances and must be considered as a mere loading parameter.. (a) Various energy contributions normalized by the thickness ($L_{z}$): external work ($W_{\rm{ext}}$), potential energy ($\Psi_{\rm{pot}}$), elastic strain energy ($\Psi_{\rm{\epsilon}}$), surface free energy ($\Psi_{\rm{SFE}}$), and other unquantified contributions (plasticity, folding, non-linear elasticity). (b) Dissipation calculated from the energy balance ($\Psi_{d}$) and crack surface energy from phase-field calculations. (c) Fracture toughness calculated using various methods.
@@
Conclusions.—This Letter sheds new light on the origin of fracture toughness in amorphous silica, an archetype of brittle materials. We investigated the underlying mechanisms that explain the difference between free surface energy and fracture toughness in amorphous silica at the crossover between the atomic and the continuum scales. Using a combination of molecular dynamics and phase-field modeling, we quantified the different contributions to the energy dissipation that occurs during fracture. From a comprehensive energy balance analysis, we were able to clearly isolate the free surface energy term and identify an additional contribution resulting from damage, not plastic deformation. This damage contribution exceeds the free surface energy by a factor of about 4, in agreement with the experimental measurements.
@@
These findings highlight that traditional notions of fracture in brittle materials may overlook the subtle nature of mechanical dissipation in fracture of some amorphous materials. In particular they emphasize the dominant contribution of diffused damage in silicate glasses, aligning with experimental observations. These results open novel perspectives beyond silicate glasses. For example, it would be interesting to evaluate if these concepts are relevant for some metallic glasses [36], which are known to span the transition from brittle to ductile fracture. How they couple to dynamic fracture is also a question worth investigating. The phase-field description is uniquely qualified to identify damage in the presence of plasticity. Therefore, it offers a more nuanced perspective for future materials design, particularly in optimizing the mechanical properties of brittle materials like amorphous silicates where, e.g., the addition of network modifiers may increase the contribution of plasticity.

### v4

@@ 3.2 Local plasticity
There are two essential phenomenological differences between damaging and ductile behavior: (i) in damage, no plateau is observed; instead, the stress reaches a maximum and then decreases to zero with further loading; (ii) the elastic stiffness in damage is lost, while in plasticity it is conserved upon unloading. These behaviors are illustrated in Fig. 3a.
Furthermore, to demonstrate that plastic effects are minimal during crack propagation, we present the von Mises stress divided by the maximum value throughout the entire loading process for a crack in Fig. 3b. It is evident that in the damaged zone, this ratio tends to zero, indicating damage. In contrast, if the material were ductile, the ratio would have remained around 1. Furthermore, this is reinforced by the energy balance presented in Section 4, which shows that the unaccounted energies are minimal.
3.3 Free surface energy
The free surface energy ($2\gamma$) was determined using a smaller cubic sample with a side length of 100 Å. We found that, the identified values to be independent of the sample size once it exceeded the 50 Å box length.
We determined $2\gamma=2.8\pm 0.2$ J/m2. This value is somewhat larger than room temperature experimental measurements including humidity ($\approx 0.5$ J/m2) [Wiederhorn69, Sarlat06, Kimura2015] but in agreement with calculated temperature variations [Yu2018] and correlates well with simulations using more sophisticated potentials [Rimsza2018].
Having access to local quantities gives us a unique opportunity to identify the origin and the distribution of the changes due to the newly formed surface.
By looking at atomistic displacements, we observe that after cutting, equilibrium is reached through atomic displacements localized within a few interatomic distances from the surface. This finding aligns well with experimental observations of surface relaxation in crystals, where an exponential decay over the first few atomic layers has been reported [Desjonqueres2012].
Secondly, the local change in potential energy can be determined using a coarse-graining technique to estimate its spatial distribution. The coarse-graining was performed on the initial coordinates to maintain continuity, and the potential energy density values shown in Fig. 4a were shifted to the center for clearer visualization.
Fig. 4a shows that as a function of coarse-graining width, the profile of $d\psi_{\rm{SFE}}$ changes and reflects the coarse-graining function. This result indicates that the energy change is localized near the surface, and this change does not diffuse into the sample. However, when integrating this change along the $y$ axis, we recover a local $2\gamma$, whose mean value and standard deviation are displayed in Fig. 4b, normalized by the global value without coarse-graining. It is important to observe that the mean value, and thus the overall sum, is independent of the coarse-graining width; only the finer details are lost when $w$ is increased. Consequently, when determining the overall change in surface energy, the coarse-graining procedure does not affect the results.
Figure 4: Distribution of free surface energy as a function of coarse-graining width. (a) Profiles of free surface energy along the $y$ axis for various coarse-graining widths. (b) Mean value and standard deviation of the local free surface energy $\gamma$, normalized by the global value.
This increase in energy is a clear signature of a newly created surface. Unfortunately, in a mechanically deformed sample, changes in potential energy are masked by elastic deformation, i.e., the strain energy. Nevertheless, for purely elastic deformation, the potential energy $\psi_{\mathrm{pot}}$ is expected to be equal to the elastic energy $\psi_{\varepsilon}$ defined in eq. 4. In crack propagation, structural evolution—typically not accessible through purely elastic deformation—is reflected in deviations of $\psi_{\mathrm{pot}}$. We therefore introduce the difference, denoted
$\psi_{\rm{FSE}}=\psi_{\rm{pot}}-\psi_{\varepsilon}.$ (10)
reflecting the isolated free surface energy.
[…]
@@ 3.3 Free surface energy
Fig. 5a shows the distribution of $\psi_{\rm{FSE}}$ along the normal to the surface for the dissected sample. After coarse-graining, the profile width of $\psi_{\rm{SFE}}$ is observed to be 8 Å, reflecting the convolution of the highly localized peak with the width $w$. A coarse-grained profile for the cracked samples, taken along the line indicated in (b), is also shown. The two coarse-grained $\psi_{\rm{FSE}}$ distributions are similar. This observation suggests that the $\psi_{\rm{FSE}}$ calculated in the fractured sample is the energy change due to the formation of the free surface. Fig. 5b depicts the free surface energy during crack propagation.
@@
Since the energy localization is consistent across different sample conditions, it appears that the free surface energy $2\gamma$ could also be an effective measure for assessing the area of newly formed cracks. Fig. 5c displays crack length calculated from the global sample free surface energy $\Psi_{\rm{FSE}}$ divided by $2\gamma$. It is compared with the crack length determined by the FEMU based on the damage approach eq. (6). This agreement between the two methods validates the use of $\psi_{\rm{FSE}}$ as a reliable metric for quantifying crack propagation, providing an alternative to more traditional methods and offering precise measurements that align well with continuum-based models.
3.4 Structural signature
Atomic-scale simulations are a powerful tool because they provide not only global responses but also allow the user to probe the local structure, offering a unique way to identify the structural signature underlying global phenomena. Nevertheless, these results should be treated as material-specific and should not be generalized without further validation. Rather, they should serve as a basis for additional analyses when investigating other materials.
In order to identify the structural rearrangements associated with crack opening, two local quantities were calculated. As expected during the opening of the network, changes were anticipated in both the ring structure of the glass and the coordination number of the network-forming silicon atoms.
To quantify changes in the ring structure, we employed the algorithm proposed by Le Roux & Jund [LeRoux2010] to identify individual rings and monitor their evolution. Initially, each atom was assigned a value of $d_{\rm{R}}=0$. If, compared to the initial configuration, a ring was no longer present in the deformed structure, all atoms belonging to that ring were assigned $d_{\rm{R}}=1$. If an atom had already been associated with a broken ring, its value remained 1. Finally, a coarse-graining procedure was applied using a Gaussian smoothing kernel with a width of $w=8$ Å, resulting in Fig. 6(a).
The change in coordination number, $\delta\mathrm{CN}$, was computed in a similar manner. First, the number of oxygen atoms within 1.7 Å of each silicon atom was counted in both the initial and deformed configurations. These local coordination values were then coarse-grained using the same width $w=8$ Å. The difference between the initial and deformed coarse-grained values yielded the $\delta\mathrm{CN}$ field, shown in Fig. 6(c).
When visualizing slices of each type of structural change, it becomes evident that ring damage exhibits a broader spatial extent, resembling phase-field damage. In contrast, changes in coordination number are more localized and narrow, akin to free surface energy distributions.
@@ 3.4 Structural signature
As a consequence, our analysis of the local structure revealed that diffuse damage is primarily associated with changes in the ring structure, whereas variations in free surface energy correlate with changes in the coordination number of silicon atoms.
[…]
4 Energy equilibrium
@@
As noted in the introduction, experimental measurements have shown that the free surface energy resulting from surface relaxation accounts for only approximately 20% of the energy required to open a crack. This section casts the free surface energy and the damage energy we have calculated at the atomic scale into a global energy balance to identify the origin of this discrepancy.
@@
Fig. 7a illustrates the various energy contributions as a function of loading. The external work, $W_{\rm{ext}}$, is partitioned into several components: the elastic strain energy, $\Psi_{\epsilon}$, and the free surface energy (FSE), $\Psi_{\rm{FSE}}$, whose sum is represented as $\Psi_{\rm{pot}}$. Additionally, there is a small non-linear contribution, $\Psi_{\rm{etc}}$, and a significant unquantified remainder. The non-linear energy contribution is determined by isolating the portion of the elastic energy density that exceeds the local potential energy, which ideally remains positive under the assumption of linear elasticity. This excess energy contribution, is then subtracted from the total elastic energy. Notably, this contribution (represented by the dashed line) is small, suggesting that energy dissipation due to shear plasticity, densification or non-linear elasticity is indeed negligible. Under tensile loading followed by fracture, these mechanisms contribute only marginally to the overall energy dissipation within the material.
@@
Finally, Fig. 7c presents the fracture toughness ($g_{c}$) as a function of loading, calculated using three distinct methods: the FEMU phase-field fit (open circles), the energy balance approach ($\Psi_{d}/\Gamma_{\rm{SFE}}$) obtained from the coarse-grained atomic scale simulations and the J integral based on the coarse-grained stress and displacement fields (red filled circles). The J-integral is calculated as described in Section B.
It is evident that the critical fracture surface energy remains consistent across the three calculation methods. The resulting value yields a ratio of $2\gamma/g_{c}\approx 23$%: this value, which is determined at the atomic scale, matches the experimental findings [Wiederhorn69, Lucas1995, King2007, Mueller2015].
@@
Figure 7: Energy balance of the cracked sample as a function of the loading parameter. We note that $K_{I}$ loses its precise physical significance as the crack tip advances and should be regarded solely as a loading parameter. (a) Various energy contributions normalized by the thickness ($L_{z}$): external work ($W_{\mathrm{ext}}$), potential energy ($\Psi_{\mathrm{pot}}$), elastic strain energy ($\Psi_{\varepsilon}$), surface free energy ($\Psi_{\mathrm{SFE}}$), and other unquantified contributions (e.g., plasticity, folding, and nonlinear elasticity). (b) Dissipation computed from the energy balance ($\Psi_{d}$) and crack surface energy obtained from phase-field calculations. (c) Fracture toughness evaluated using different methods. The blue shaded region indicates the range of experimental measurements reported in Refs. [Wiederhorn69, Lucas1995, King2007, Mueller2015].
5 Conclusion
@@
This paper provides new insight into the origin of fracture toughness in amorphous silica, an archetype of brittle materials. We investigated the mechanisms responsible for the difference between free surface energy and fracture toughness in amorphous silica at the crossover between the atomic and the continuum scales. Using a combination of atomic scale simulations and phase-field modeling, we quantified the different contributions to energy dissipation during fracture. From a comprehensive energy balance analysis, we were able to clearly isolate the free surface energy term and identify an additional contribution. Our analysis shows that this contribution arises almost exclusively from damage rather than plastic deformation, contradicting the conventional assumption [Wiederhorn69]. Notably, our calculations predict that this damage contribution exceeds the free surface energy by approximately a factor of 4, in agreement with experimental measurements.
@@
These findings highlight that traditional views of fracture in brittle materials may overlook the subtle nature of mechanical dissipation in fracture of some amorphous materials. In particular they emphasize the dominant contribution of diffused damage in silicate glasses, aligning with experimental observations. These results also open new perspectives beyond silicate glasses. For example, it would be interesting to assess if these concepts apply to metallic glasses [lewandowski2005intrinsic] which are known to span the transition from brittle to ductile fracture. Understanding how these mechanisms couple to fatigue or to dynamic fracture is also a question worth investigating. The phase-field description is uniquely qualified to identify damage in the presence of plasticity. Therefore, it offers a more nuanced perspective for future materials design, particularly in optimizing the mechanical properties of brittle materials like amorphous silicates where, e.g., the addition of network modifiers may increase the contribution of plasticity.
Acknowledgment
@@
Appendix A Effect of convolution length
One of the main findings of this paper is the identification of an emerging length scale from molecular simulations using a Gaussian convolution. This homogenization process starts with a finite length, making it crucial to demonstrate that the results presented are essentially independent of the specific parameters used in the homogenization.
In this section, we explore how the width of the coarse-graining affects the results discussed in the main body of the article.
There are three key aspects that may be influenced by the homogenization process: (i) the elastic behavior, which indicates at what scale discrete atoms can be approximated as a continuum; (ii) the diffusion width of damage; and (iii) the effect of the free surface on local potential energy.
Our findings indicate that for a minimal length scale of $w=8$ Å, the elastic strain energy matches the local potential energy. This result aligns well with our previous findings on local elasticity. Additionally, we show that a width of $w=8$ Å has minimal impact on damage diffusion. Lastly, we demonstrate that the free surface energy is highly localized on the surface, implying that while varying $w$ may alter the local maximum amplitude, it does not affect the overall sum of the energy change. Detailed explanations are provided below.
Figure 8: Impact of coarse-graining width on local elastic and potential energy densities in the material’s elastic state.
Fig. 8 depicts the local strain energy density ($d\psi_{\rm{el}}$) and potential energy density ($d\psi_{\rm{pot}}$) in black and red, respectively. The strain energy density is calculated from the stresses and strains, while the potential energy is coarse-grained from the atomistic values. These values are shown as a function of the global loading at a position in the sample that remains elastic during deformation. The descent of the curves indicates unloading, not damage nor plasticity. Under elastic conditions, these two values are expected to match. However, as illustrated in Fig. 8, if the homogenization length is too small, neither quantity is well-defined. When calculating the surface free energy ($d\psi_{\rm{SFE}}=d\psi_{\rm{pot}}-d\psi_{\rm{el}}$), this can result in a negative value, particularly in regions of the model where it should be zero, which is clearly not possible. Therefore, elastic analysis provides a lower bound for the coarse-graining width, which is in the range of w=68w=6–8 Å. We note that a similar length scale, $w=8$ Å, was found where local elasticity begins to be well-defined [Molnar2016JNCS]. This was explained by micro-irreversible rearrangements present in amorphous materials even in the initial elastic stage, which are included in the slope of the experimentally measured stiffness.
[…]
It is clearly visible that for w=35w=3–5 Å, the mean value remains the same, with only the fluctuation decreasing. At $w=8$ Å, the mean value starts to increase slightly, but this increase remains within the order of 10 %. Beyond this, as $w$ increases, the width of the damage zone, $l$, increases linearly.
To keep negative surface energies minimal and stresses well-defined, we accepted a slight effect on the damage diffusion and chose $w=8$ Å for subsequent analysis. It is important to note that the exact value of $l$ might be slightly smaller. This consideration is crucial as the exact value depends on the degradation function and the phase-field description.
Appendix B J integral
The J integral, a contour integral, was originally proposed to deduce the energy liberated from an elastic body upon the potential advancement of a sharp crack within a homogeneous and isotropic domain. Essentially, the contour integral provides the difference between the work done on the contour and the energy stored in the solid. The resulting difference, assuming a unitary, straight, horizontal crack advancement, gives the energy release rate. Its advantage at the time was that it captured the additional energy dissipated by ductile deformation and attributed it to the increment of fracture toughness. Therefore, the method is applicable in ductile cases.
In this paper, we use the J integral to gain an approximate description and to verify the critical fracture toughness identified by the phase-field damage model.
Figure 10: (a) Schematic illustration to calculate the J integral. (b) Characteristic half contour width as a function of loading.
The contour integral is defined as:
$J=\int\limits_{{\Gamma_{J}}}{\left({{\psi_{el}}{n_{x}}-{\bf{t}}\frac{{\partial{\bf{u}}}}{{\partial x}}}\right)}d\Gamma$ (11)
where ${\bf{t}}={\bf{Pn}}$ with $\bf{P}$ being the coarse-grained first Piola-Kirchhoff stress tensor and $\bf{n}$ the normal vector to the contour.
The contour integral was centered at the local maximum of the first principal stress peak in front of the crack. To approximate the effect of crack undulation multiple contours were used, with widths denoted as $L_{\Gamma}$. Finally, a sigmoidal fit, which transformed into a linear function, was applied:
$J\left({{L_{\Gamma}}}\right)=\frac{{{J_{\infty}}}}{{1+{e^{-b{L_{\Gamma}}}}}}+{L_{\Gamma}}c-\frac{{{J_{\infty}}}}{2}.$ (12)
This fitting procedure is displayed in Fig. 10a.
Due to local damage and the inhomogeneous nature of the material, this contour integral is not well-defined at the crack tip. Additionally, the function exhibits a slight slope at larger contour sizes because the expected K-field is influenced by the model’s boundary conditions. Therefore, the results shown in Section 4 display $J_{0}$, which is a linear interpolation of $J$ at $L_{\Gamma}=0$. Interestingly, this fit allows us to identify the length where the integral is valid and the response behaves closely to linear elastic fracture mechanics. This is done by intersecting the initial slope and the line from the final slope of eq. (12), as shown in Fig. 10a. The length $l_{J}$ is displayed in Fig. 10b. It shows a similar size to where the material becomes treatable with linear elastic fracture mechanics when the contour becomes larger than the size of the diffused damaged zone.